# On semidefinite descriptions for convex hulls of quadratic programs

### 07 Mar 2024 v1

@@ B.1 Mixed binary programming
${\cal R}^{\prime}(x,t)$ =$=\left\{\begin{array}[]{r}\begin{pmatrix}2t\\ -x_{2}\\ 0\end{pmatrix},\,\begin{pmatrix}-x_{2}\\ x_{1}\\ 0\end{pmatrix},\,\begin{pmatrix}t\\ -x_{2}\\ x_{1}\end{pmatrix},\,\begin{pmatrix}x_{2}(x_{1}-1+2t)\\ -x_{1}^{2}+x_{1}-2tx_{1}-2t\\ 2x_{2}\end{pmatrix}\end{array}\right\}^{\perp}.$ (20)
In this case, one may verify that $(x,t)\in{\cal R}^{\prime}(x,t)$. Then, as $(x,t)$ is nonzero we deduce that ${\cal R}^{\prime}(x,t)$ is nontrivial in this case as well.
As ${\cal R}^{\prime}(x,t)$ is nontrivial in all three cases, we conclude from Theorem 2 that
$\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}=\left\{\begin{array}[]{l}(x,t)\in{\mathbb{R}}^{3}:\\ \leavevmode\nobreak\ \leavevmode\nobreak\ -(q_{\textup{obj}}(x)-2t)-q_{1}(x)\\ \qquad\geq\sqrt{(q_{\textup{obj}}(x)-2t-q_{1}(x))^{2}+2q_{2}(x)^{2}}\end{array}\right\}.$
Here, the second equation follows from the last equivalence of ${\cal S}_{\textup{SDP}}$ in Remark 1 and the characterization of $\Gamma^{\circ}$ derived as the rotated second-order cone for this specific set ${\cal S}$. Substituting back the definitions of $q_{\textup{obj}},q_{1},q_{2}$ and squaring both sides of the constraint in the set ${\cal S}_{\textup{SDP}}$ above gives
$\operatorname{conv}({\cal S})=\left\{(x,t)\in{\mathbb{R}}^{3}:\,\begin{array}[]{l}-(x_{2}^{2}-2t)-x_{1}(x_{1}-1)\geq 0\\ 4(1-x_{1})(2tx_{1}-x_{2}^{2})\geq 0\end{array}\right\}.$
While the two descriptions of $\operatorname{conv}({\cal S})$ given in (6) and given above are different, we can check that they indeed define the same set. First, the $\subseteq$ inclusion follows immediately as $x_{2}^{2}-2t\leq x^{2}-2tx_{1}\leq 0$ so that both constraints in the description above are satisfied. In the reverse direction, we first establish that every $(x,t)$ in the set above satisfies $x_{1}\in[0,1]$. If $x_{1}<0$, then the first constraint implies $t>0$ while the second constraint implies $t\leq 0$, a contradiction. If $x_{1}>1$, then the second constraint implies $2t\leq x_{2}^{2}/x_{1}$. On the other hand, the first constraint implies $0<x_{1}(1-x_{1})\leq-(x_{2}^{2}-2t)\leq x_{2}^{2}(1/x_{1}-1)$, which contradicts $x_{1}>1$. Finally, suppose $(x,t)$ is in the set above with $x_{1}\in[0,1]$. We will show that $(x,t)$ satisfies the constraints in (6). If $x_{1}=1$, then $x_{2}^{2}-2tx_{1}=x_{2}^{2}-2t\leq 0$ by the first constraint (whence also $t\geq 0$). If $x_{1}\in(0,1)$, then the second constraint is equivalent to $x_{2}^{2}-2tx_{1}\leq 0$ (again, $t\geq 0$). Finally, if $x_{1}=0$, then we have $x_{2}^{2}-2tx_{1}\leq 0$ by the second constraint and $x_{2}^{2}-2t\leq 0$ by the first constraint (so that $t\geq 0$).
B.2 Quadratic matrix programs
@@
Quadratic matrix programs (QMPs) [2, 25] are a generalization of QCQPs where the decision variable $x\in{\mathbb{R}}^{n}$ is replaced by a decision matrix $X\in{\mathbb{R}}^{r\times k}$. These problems find a variety of application and have been used to model robust least squares problems, the orthogonal Procrustes problem, and certain sphere packing problems. Formally, a QMP is an optimization problem in the variable $X\in{\mathbb{R}}^{r\times k}$, where the constraints and objective function are each of the form
@@
We will choose to view QMPs as the special class of QCQPs where the quadratic forms $A_{\textup{obj}},A_{1},\dots,A_{m}$ are each of the form $I_{k}\otimes\mathbb{A}$ for some $\mathbb{A}\in{\mathbb{S}}^{r}$.
We will next establish that if the number of constraints is small compared to $k$ (originally the width of the matrix variable), then convex hull exactness holds. Specifically, we show below that if Assumption 1 holds, $A_{\textup{obj}}=I_{k}\otimes\mathbb{A}_{\textup{obj}}$, $A_{1}=I_{k}\otimes\mathbb{A}_{1}$, …, $A_{m}=I_{k}\otimes\mathbb{A}_{m}$ for some $\mathbb{A}_{\textup{obj}},\mathbb{A}_{1},\dots,\mathbb{A}_{m}\in{\mathbb{S}}^{r}$ and $k\geq m$, then ${\cal R}^{\prime}(x,t)$ is nontrivial for every $(x,t)\in{\cal S}_{\textup{SDP}}\setminus{\cal S}$. By Theorem 1, this will imply that convex hull exactness holds, i.e., $\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}$.
@@
${\cal R}^{\prime}(x,t)=\left\{(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}:\,\begin{array}[]{l}x^{\prime\top}A[\gamma]x^{\prime}=0,\,\forall(1,\gamma)\in{\cal G}(x,t)^{\perp}\\ \left\langle A[\gamma]x+b[\gamma],x^{\prime}\right\rangle-t^{\prime}=0,\,\forall(1,\gamma)\in{\cal G}(x,t)^{\perp}\end{array}\right\}$ (23)
@@
$\left\langle\begin{pmatrix}\gamma_{\textup{obj}}\\ \gamma\end{pmatrix},q(x)-2te_{\textup{obj}}+\begin{pmatrix}\ell_{\textup{obj}}\\ \ell\end{pmatrix}\right\rangle\leq\left\langle\begin{pmatrix}\gamma_{\textup{obj}}\\ \gamma\end{pmatrix},q(x)-2te_{\textup{obj}}+\begin{pmatrix}\left\langle A_{\textup{obj}},Y\right\rangle\\ \left(\left\langle A_{i},Y\right\rangle\right)_{i}\end{pmatrix}\right\rangle\leq 0,$
where the last inequality follows from (25) and $(\gamma_{\textup{obj}},\gamma)\in\Gamma\subseteq{\mathbb{R}}^{1+m}_{+}$. This then shows that $q(x)-2te_{\textup{obj}}+(\ell_{\textup{obj}},\ell)\in\Gamma^{\circ}$. Moreover, because $\mathbb{A}(\gamma_{\textup{obj}},\gamma)\succeq 0$, we have
@@
SDP exactness in the context of QMPs was previously studied by Wang and Kılınç-Karzan [25], Beck [2]. Specifically, under Assumption 1, Beck [2] shows that objective value exactness holds whenever $k\geq m$ and Wang and Kılınç-Karzan [25] show that convex hull exactness holds whenever $k\geq m+2$. The above calculation strengthens both of these results by showing that under Assumption 1 convex hull exactness holds whenever $k\geq m$.
B.3 Sets defined by two quadratic constraints
We apply our results to derive a characterization of convex hull exactness of SDP relaxation when we have only two quadratic constraints. Note that when such a set is obtained from the epigraphical representation of a quadratic objective function and a single quadratic constraint, SDP convex hull exactness is well-studied, see e.g., [24] and references therein. In contrast, we will show below that the case of two general quadratic constraints is more nuanced. In the case of strict inequalities and under a strict feasibility assumption, Yıldıran [26] establishes that the convex hull of a set defined by two quadratic constraints is given by two constraints admitting second-order cone (SOC) representations and these constraints can be obtained by examining the eigenvalues of certain Lagrange aggregation matrices. Our results will present a simple way of checking exactly when such a convex hull characterization is given by its SDP relaxation as opposed to the more involved condition of [26].
Consider a set defined by two quadratic constraints, $\hat{\cal S}\coloneqq\left\{x\in{\mathbb{R}}^{n}:\,q_{1}(x)\leq 0,q_{2}(x)\leq 0\right\}$. To connect such a set to the epigraphical sets studied in the main body, we will define $q_{\textup{obj}}(x)\coloneqq 0$. Then, Assumption 1 is equivalent to the assumption that there exists $\hat{\gamma}=(\hat{\gamma}_{1},\hat{\gamma}_{2})\in{\mathbb{R}}^{2}_{+}$ such that $\hat{\gamma}_{1}A_{1}+\hat{\gamma}_{2}A_{2}\succ 0$. The projected SDP relaxation of this set is given by
@@ B.3 Sets defined by two quadratic constraints
Note that $\Gamma_{1}$ is a cone in ${\mathbb{R}}^{2}$ so that it is necessarily polyhedral. We deduce that $\Gamma$ is also polyhedral and we may apply Section 3.4 in this situation. Note that as $A[\hat{\gamma}]\succ 0$, we must have that $\Gamma_{1}$ is full dimensional. Let $\gamma^{(1)},\gamma^{(2)}\in{\mathbb{R}}^{2}_{+}$ be generators of $\Gamma_{1}$, i.e., such that $\Gamma_{1}=\operatorname{cone}(\left\{\gamma^{(1)},\gamma^{(2)}\right\})$. Then, $\Gamma=\operatorname{cone}\left(\left\{\begin{pmatrix}0\\ \gamma^{(1)}\end{pmatrix},\begin{pmatrix}0\\ \gamma^{(2)}\end{pmatrix},e_{\textup{obj}}\right\}\right)$ so that $\Gamma$ has exactly 6 proper faces that we deal with individually below.
@@
$\begin{cases}[\gamma^{(1)},q(x)]=0,\\ [\gamma^{(2)},q(x)]<0.\end{cases}$
@@
We collect the above discussion in the following lemma.
@@
In particular, whenever we have $q_{1}$ and $q_{2}$ are both nonconvex, primal strict feasibility is satisfied (i.e., $\exists\hat{x}\in{\mathbb{R}}^{n}$ such that $q_{1}(\hat{x}),q_{2}(\hat{x})<0$) and dual strict feasibility is satisfied (i.e., $\exists\hat{\gamma}\in{\mathbb{R}}^{2}_{+}$ such that $A[\hat{\gamma}]\succ 0$), then $\operatorname{conv}(\hat{\cal S})=\hat{\cal S}_{\textup{SDP}}$ if and only if for both $i=1,2$, we have that
@@
Lemma 4 presents an easy to check condition for SDP convex hull exactness in the case of two quadratic constraints. One may compare this to the main result in [26] which gives a description of $\operatorname{clconv}(\hat{\cal S})$ assuming that $\hat{\cal S}$ is defined by strict inequality constraints.
We explain the description of [26] using a mix of our notation and the notation in [26]. Let
$Q_{i}\coloneqq\left(\begin{smallmatrix}A_{i}&b_{i}\\ b_{i}^{\intercal}&c_{i}\end{smallmatrix}\right),\qquad\forall i=1,2.$
For any $\gamma\in{\mathbb{R}}^{2}$, define $Q[\gamma]\coloneqq\gamma_{1}Q_{1}+\gamma_{2}Q_{2}$. Let $\lambda_{i}(\cdot)$ denote the $i$th smallest eigenvalue of its argument. Then, under the assumption that $\hat{\cal S}$ is defined via strict inequalities, is nonempty, and $\operatorname{clconv}(\hat{\cal S})\neq{\mathbb{R}}^{n}$, [26] establishes that
$\operatorname{clconv}(\hat{\cal S})=\left\{x\in{\mathbb{R}}^{n}:\,[\gamma_{Y}^{(1)},q(x)]\leq 0,\,[\gamma_{Y}^{(2)},q(x)\leq 0]\right\},$
where $\gamma_{Y}^{(1)},\gamma_{Y}^{(2)}\in{\mathbb{R}}^{2}_{+}$ are vectors that generate the connected component (a cone) of
$\left\{\gamma\in{\mathbb{R}}^{2}_{+}:\,\lambda_{2}(Q[\gamma])\geq 0\right\}$
containing $\Gamma_{1}$. By the definition of $\gamma_{Y}^{(i)}$, the set $\left\{x\in{\mathbb{R}}^{n}:[\gamma_{Y}^{(i)},q(x)]\leq 0\right\}$ is either convex (and thus SOC representable in the homogenized space) or it is the union of an SOC and its negative (in the homogenized space) referred as a semi-cone. In the latter case, [26] further establishes that we can in fact restrict our attention to only one side of the associated nonconvex semi-cone to get a convex representation of $\operatorname{clconv}(\hat{\cal S})$.
To summarize, in order to test SDP exactness based on the characterization of [26], we will have to first compute $\operatorname{clconv}(\hat{\cal S})$ by computing $\gamma_{Y}^{(1)},\gamma_{Y}^{(2)}\in{\mathbb{R}}^{2}_{+}$ from $\left\{\gamma\in{\mathbb{R}}^{2}_{+}:\,\lambda_{2}(Q[\gamma])\geq 0\right\}$ and then check whether the set $\left\{x\in{\mathbb{R}}^{n}:\,[\gamma_{Y}^{(i)},q(x)]\leq 0,\,\forall i\in[2]\right\}$ is the same as the one given by $\left\{x\in{\mathbb{R}}^{n}:\,[\gamma^{(i)},q(x)]\leq 0,\,\forall i\in[2]\right\}$ obtained from $\Gamma_{1}$. In contrast to this our condition given in Lemma 4 simply requires us to compute $\Gamma_{1}=\operatorname{cone}\left\{\gamma^{(1)},\gamma^{(2)}\right\}$ and check for both $i\in[2]$ whether $\ker(A[\gamma^{(i)}])\cap b[\gamma^{(i)}]^{\perp}$ is nontrivial whenever $\gamma^{(i)}\in{\mathbb{R}}^{2}_{++}$; this can be done by solving a system of linear equations. In the rest of this remark we establish the equivalence of our condition and the one from [26] when SDP exactness holds.
To explain the presence of the second eigenvalue of $Q[\gamma]$, note that by assumption (of strict feasibility of $\hat{\cal S}$), there exists an $\hat{x}\in{\mathbb{R}}^{n}$ for which $q_{1}(\hat{x}),q_{2}(\hat{x})<0$ so that for any nonzero $\gamma\in{\mathbb{R}}^{2}_{+}$, we have
$\begin{pmatrix}\hat{x}\\ 1\end{pmatrix}^{\intercal}Q[\gamma]\begin{pmatrix}\hat{x}\\ 1\end{pmatrix}<0.$
We deduce that $Q[\gamma]$ has at least one negative eigenvalue for all nonzero $\gamma\in{\mathbb{R}}^{2}_{+}$. On the other hand, for any $\gamma\in\Gamma_{1}$, by Cauchy’s Eigenvalue Interlacing Theorem, we have that
$\lambda_{2}(Q[\gamma])\geq\lambda_{1}(A[\gamma])\geq 0$
so that $Q[\gamma]$ has exactly one negative eigenvalue for all nonzero $\gamma\in\Gamma_{1}$.
Now, outside of a few special edge cases (e.g., when $\operatorname{clconv}(\hat{\cal S})$ is defined by a single convex constraint), $\operatorname{clconv}(\hat{\cal S})=\hat{\cal S}_{\textup{SDP}}$ if and only if $\Gamma_{1}=\operatorname{cone}\left\{\gamma_{Y}^{(1)},\gamma_{Y}^{(2)}\right\}$, i.e., $\Gamma_{1}$ coincides with one of the connected components of $\left\{\gamma\in{\mathbb{R}}^{2}_{+}:\,\lambda_{2}(Q[\gamma])\geq 0\right\}$. The latter condition holds if and only if for both $i=1,2$, either $\gamma^{(i)}$ lies on the boundary of +${\mathbb{R}}^{2}_{+}$ or that $\lambda_{2}(Q[\gamma^{(i)}-\epsilon\hat{\gamma}])<0$ for all $\epsilon>0$ small enough, i.e., that the number of negative eigenvalues in $Q$ increases when moving out of $\Gamma_{1}$.
The remainder of this remark shows that this condition is equivalent to our requirement that $\ker(A[\gamma^{(i)}])\cap b[\gamma^{(i)}]^{\perp}$ is nontrivial if $\gamma^{(i)}\in{\mathbb{R}}^{2}_{++}$. To this end, suppose $\gamma^{(1)}\in{\mathbb{R}}^{2}_{++}$. The reasoning for $\gamma^{(2)}\in{\mathbb{R}}^{2}_{++}$ is analogous.
First, suppose that $\ker(A[\gamma^{(1)}])\cap b(\gamma^{(1)})^{\perp}$ is nontrivial and let $\xi$ be a unit vector from $\ker(A[\gamma^{(1)}])\cap b(\gamma^{(1)})^{\perp}$. Then, as $\gamma^{(1)}\in\Gamma_{1}$ is nonzero we deduce that $Q[\gamma^{(1)}]$ has exactly one negative eigenvalue. Furthermore,
$Q[\gamma^{(1)}]\begin{pmatrix}\xi\\ 0\end{pmatrix}=\begin{pmatrix}A[\gamma^{(1)}]\xi\\ b[\gamma^{(1)}]^{\intercal}\xi\end{pmatrix}=0$
so that $(\xi,0)$ is in $\ker(Q[\gamma^{(1)}])$ and hence $\lambda_{2}(Q[\gamma^{(1)}])=0$. Let $(\eta,t)$ denote a unit eigenvector corresponding to $\lambda_{1}(Q[\gamma^{(1)}])$. Our goal is to show that $\lambda_{2}(Q[\gamma^{(1)}-\epsilon\hat{\gamma}])<0$ for all $\epsilon>0$ small enough. By the variational characterization of $\lambda_{2}$, it suffices to show that the quadratic form induced by $Q[\gamma^{(1)}-\epsilon\hat{\gamma}]$ is negative for all $\epsilon>0$ small enough on some two-dimensional subspace (other than at zero). We will use the subspace spanned by $(\xi,0)$ and $(\eta,t)$ for this purpose. Let ${\mathbf{S}}^{1}\subseteq{\mathbb{R}}^{2}$ denote the unit circle and let $\epsilon>0$. Abusing notation, we will identify $v\in{\mathbf{S}}^{1}\subseteq{\mathbb{R}}^{2}$ with the unit vector $v_{1}(\xi;0)+v_{2}(\eta;t)$. Then,
$\lambda_{2}(Q[\gamma^{(1)}-\epsilon\hat{\gamma}])$ $\leq\max_{v\in{\mathbf{S}}^{1}}\left\{v^{\intercal}Q[\gamma^{(1)}-\epsilon\hat{\gamma}]v\right\}$
$=\max_{v\in{\mathbf{S}}^{1}}\left\{\begin{array}[]{l}\left(v_{1}^{2}\xi^{\intercal}A[\gamma^{(1)}-\epsilon\hat{\gamma}]\xi\right)\\ \quad+2v_{1}v_{2}\begin{pmatrix}\xi\\ 0\end{pmatrix}^{\intercal}Q[\gamma^{(1)}-\epsilon\hat{\gamma}]\begin{pmatrix}\eta\\ t\end{pmatrix}\\ \quad+v_{2}^{2}\begin{pmatrix}\eta\\ t\end{pmatrix}^{\intercal}Q[\gamma^{(1)}-\epsilon\hat{\gamma}]\begin{pmatrix}\eta\\ t\end{pmatrix}\end{array}\right\}$
$=\max_{v\in{\mathbf{S}}^{1}}\left\{\begin{array}[]{l}-\epsilon v_{1}^{2}\xi^{\intercal}A[\hat{\gamma}]\xi-2\epsilon v_{1}v_{2}\begin{pmatrix}\xi\\ 0\end{pmatrix}^{\intercal}Q[\hat{\gamma}]\begin{pmatrix}\eta\\ t\end{pmatrix}\\ \quad+v_{2}^{2}\lambda_{1}(Q[\gamma^{(1)}])-\epsilon v_{2}^{2}\begin{pmatrix}\eta\\ t\end{pmatrix}^{\intercal}Q[\hat{\gamma}]\begin{pmatrix}\eta\\ t\end{pmatrix}\end{array}\right\}$
$=\max_{v\in{\mathbf{S}}^{1}}v^{\intercal}\begin{pmatrix}-\epsilon\xi^{\intercal}A[\hat{\gamma}]\xi&\Theta(\epsilon)\\ \Theta(\epsilon)&\lambda_{1}(Q[\gamma^{(1)}])+\Theta(\epsilon)\end{pmatrix}v.$
Here, the $\Theta(\epsilon)$ denote functions of $\epsilon$ with magnitude bounded by $O(\epsilon)$. Thus, as $A[\hat{\gamma}]\succ 0$ and $\lambda_{1}(Q[\gamma^{(1)}])<0$, we deduce that the two-by-two matrix in the final line is negative definite so that $\lambda_{2}(Q[\gamma^{(1)}-\epsilon\hat{\gamma}])<0$ for all $\epsilon>0$ small enough.
In the reverse direction, suppose $\lambda_{2}(Q[\gamma^{(1)}-\epsilon\hat{\gamma}])<0$ for all $\epsilon>0$ small enough. By continuity of eigenvalues and the fact that $\lambda_{2}(Q[\gamma^{(1)}])\geq\lambda_{1}(A[\gamma^{(1)}])=0$, it must hold that $\lambda_{2}(Q[\gamma^{(1)}])=0$. We may then invoke the equality case of Cauchy’s Eigenvalue Interlacing Theorem [13, Theorem 4.3.17] to conclude that $\ker(A[\gamma^{(1)}])\cap b[\gamma^{(1)}]^{\perp}$ is nontrivial.

### 20 Mar 2024 v2

@@ 3.4 Revisiting the setting of polyhedral $\Gamma$
${\cal R}^{\prime}(x,t)=\left\{(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}:\,\begin{array}[]{l}x^{\prime}\in\ker(A[f])\\ \left\langle b[\gamma],x^{\prime}\right\rangle-t^{\prime}=0,\,\forall(1,\gamma)\in{\cal F}(x,t)\end{array}\right\}.$
The only dependence on $(x,t)$ in the above description of ${\cal R}^{\prime}$ is in the definition of ${\cal F}={\cal F}(x,t)$. Hence, $\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}$ if and only if the above set is nontrivial for every ${\cal F}\trianglelefteq\Gamma$ which is exposed by some vector $q(x)-2te_{\textup{obj}}$ for $(x,t)\in{\cal S}_{\textup{SDP}}\setminus{\cal S}$.
Proof.
We begin by noting that when $\Gamma$ is polyhedral, so too is $\Gamma^{\circ}$ so that Assumption 2 holds. Next, we claim that for every face ${\cal G}\trianglelefteq\Gamma^{\circ}$ we have ${\cal G}^{\perp}=\operatorname{span}({\cal G}^{\triangle})$. By definition, $\operatorname{span}({\cal G}^{\triangle})=\operatorname{span}(\Gamma\cap{\cal G}^{\perp})\subseteq{\cal G}^{\perp}$. On the other hand, as $\Gamma$ and $\Gamma^{\circ}$ are polyhedral, we have that $\dim({\cal G})+\dim({\cal G}^{\triangle})=m$ [17, Theorem 3]. Rearranging this equation, we have $\dim({\cal G}^{\triangle})=m-\dim({\cal G})=\dim({\cal G}^{\perp})$. We conclude that ${\cal G}^{\perp}=\operatorname{span}({\cal G}^{\triangle})$.
Here, the second line follows as ${\cal G}^{\perp}=\operatorname{span}({\cal G}^{\triangle})$ holds for every face ${\cal G}\trianglelefteq\Gamma^{\circ}$ and by definition ${\cal F}(x,t)={\cal G}(x,t)^{\triangle}$. The third line follows from the fact that $(1,f)\in\operatorname{rint}({\cal F}(x,t))$ implies $\ker(A[f])\subseteq\ker(A[\gamma])$ for every $(1,\gamma)\in{\cal F}(x,t)$. The result then follows from theorem 2. ∎
Remark 4.
The main difference between section 3.4 and [18, Theorem 1] is that section 3.4 considers certain (a fortiori semidefinite) faces of $\Gamma$ whereas [18, Theorem 1] imposes a constraint on every semidefinite face of $\Gamma$. This idea of restricting the analysis to certain faces of $\Gamma$ was previously investigated by [6, 12] who used it to provide sufficient conditions for objective value exactness.
4 Example applications of theorems 1, 2 and 3.4
We apply the results of section 3 to two examples.
4.1 Quadratic matrix programs
@@
Quadratic matrix programs (QMPs) [2, 18] are a generalization of QCQPs where the decision variable $x\in{\mathbb{R}}^{n}$ is replaced by a decision matrix $X\in{\mathbb{R}}^{r\times k}$. These problems find a variety of applications, e.g., robust least squares problems and sphere packing problems. Formally, a QMP is an optimization problem in the variable $X\in{\mathbb{R}}^{r\times k}$, where the constraints and objective function are each of the form
@@
Thus, we can view QMPs as the special class of QCQPs where $A_{\textup{obj}},A_{1},\dots,A_{m}$ are each of the form $I_{k}\otimes\mathbb{A}$ for some $\mathbb{A}\in{\mathbb{S}}^{r}$.
Below, we establish a sufficient condition for convex hull exactness in this setting. Specifically, we will show that if Assumption 1 holds, for each $i\in[0,m]$, we can write $A_{i}=I_{k}\otimes\mathbb{A}_{i}$ for some $\mathbb{A}_{i}\in{\mathbb{S}}^{r}$, and $k\geq m$, then ${\cal R}^{\prime}(x,t)$ is nontrivial for every $(x,t)\in{\cal S}_{\textup{SDP}}\setminus{\cal S}$. By theorem 1, this implies that $\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}$.
@@
${\cal R}^{\prime}(x,t)$ =$=\left\{(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}:\,\begin{array}[]{l}\quad x^{\prime\top}A[\gamma]x^{\prime}=0,\,\forall(1,\gamma)\in{\cal G}(x,t)^{\perp}\\ \quad\left\langle A[\gamma]x+b[\gamma],x^{\prime}\right\rangle-t^{\prime}=0,\,\forall(1,\gamma)\in{\cal G}(x,t)^{\perp}\end{array}\right\}$ (9)
@@
$\left\langle\begin{pmatrix}\gamma_{\textup{obj}}\\ \gamma\end{pmatrix},q(x)-2te_{\textup{obj}}+\begin{pmatrix}\ell_{\textup{obj}}\\ \ell\end{pmatrix}\right\rangle$
$\leq\left\langle\begin{pmatrix}\gamma_{\textup{obj}}\\ \gamma\end{pmatrix},q(x)-2te_{\textup{obj}}+\begin{pmatrix}\left\langle A_{\textup{obj}},Y\right\rangle\\ \left(\left\langle A_{i},Y\right\rangle\right)_{i}\end{pmatrix}\right\rangle\leq 0,$
where the last inequality follows from (11) and $(\gamma_{\textup{obj}},\gamma)\in\Gamma\subseteq{\mathbb{R}}^{1+m}_{+}$. Thus, $q(x)-2te_{\textup{obj}}+(\ell_{\textup{obj}},\ell)\in\Gamma^{\circ}$. Moreover, as $\mathbb{A}(\gamma_{\textup{obj}},\gamma)\succeq 0$, we have
@@
SDP exactness for QMPs was previously studied by Wang and Kılınç-Karzan [18], Beck [2]. Specifically, under Assumption 1, Beck [2] shows that objective value exactness holds if $k\geq m$ and Wang and Kılınç-Karzan [18] show that convex hull exactness holds if $k\geq m+2$. The above calculation strengthens both of these results by showing that under Assumption 1 convex hull exactness holds if $k\geq m$.
4.2 Sets defined by two quadratic constraints
Next, consider convex hull exactness for sets defined by two quadratic constraints. In the case of strict inequalities and under a strict feasibility assumption, Yıldıran [19] establishes that the convex hull of such a set is given by two second-order cone constraints that can be obtained from the eigenvalues of an associated matrix pencil. Below, we will characterize when this convex hull is equal to its standard SDP relaxation.
Consider a set defined by two quadratic constraints, $\hat{\cal S}\coloneqq\left\{x\in{\mathbb{R}}^{n}:\,q_{1}(x)\leq 0,q_{2}(x)\leq 0\right\}$. To connect such a set to the epigraphical sets studied previously in this paper, we will define $q_{\textup{obj}}(x)\coloneqq 0$. Then, Assumption 1 states that there exists $\hat{\gamma}=(\hat{\gamma}_{1},\hat{\gamma}_{2})\in{\mathbb{R}}^{2}_{+}$ such that $\hat{\gamma}_{1}A_{1}+\hat{\gamma}_{2}A_{2}\succ 0$. The projected SDP relaxation of this set is given by
@@ 4.2 Sets defined by two quadratic constraints
Note that $\Gamma_{1}$ is a cone in ${\mathbb{R}}^{2}$ so that it is necessarily polyhedral. We deduce that $\Gamma$ is also polyhedral and we may apply section 3.4. As $A[\hat{\gamma}]\succ 0$, $\Gamma_{1}$ is full dimensional. Let $\gamma^{(1)},\gamma^{(2)}\in{\mathbb{R}}^{2}_{+}$ satisfy $\Gamma_{1}=\operatorname{cone}(\left\{\gamma^{(1)},\gamma^{(2)}\right\})$. Then, $\Gamma=\operatorname{cone}\left(\left\{(0,\gamma^{(1)}),(0,\gamma^{(2)}),e_{\textup{obj}}\right\}\right)$ so that $\Gamma$ has exactly 6 proper faces that we deal with individually below.
@@
$[\gamma^{(1)},q(x)]=0\quad\text{and}\quad[\gamma^{(2)},q(x)]<0.$
@@
We summarize this discussion below.
@@
In particular, if $q_{1}$ and $q_{2}$ are both nonconvex, primal and dual strict feasibility are satisfied (i.e., $\exists\hat{x}\in{\mathbb{R}}^{n}$ such that $q_{1}(\hat{x}),q_{2}(\hat{x})<0$ and $\exists\hat{\gamma}\in{\mathbb{R}}^{2}_{+}$ such that $A[\hat{\gamma}]\succ 0$), then $\operatorname{conv}(\hat{\cal S})=\hat{\cal S}_{\textup{SDP}}$ if and only if for both $i=1,2$, we have that
@@
Lemma 4 presents an easy to check condition for SDP convex hull exactness in the case of two quadratic constraints. In contrast, testing SDP exactness via the main result of [19] requires first the nontrivial task of computing $\operatorname{clconv}(\hat{\cal S})$ and then checking if it is equal to $\hat{\cal S}_{\textup{SDP}}$.
Acknowledgments
This work was supported by NSF [CMMI 1454548], ONR [N00014-19-1-2321], and AFOSR [FA95502210365].